OpenStudy (anonymous):

Eric is 6 1/6 feet tall and his brother is 5 3/4 feet tall. Eric is how many feet taller than his brother?

OpenStudy (hlares):

Like with the plant problem, you will first want to convert the heights of Eric and his brother intro fractions, make sure they have the same denominator, then subtract to find the difference. First, converting the heights into fractions: For Eric's height of 6 and 1/6 feet, you will want to turn 6 into a fraction with the same denominator as 1/6, then add them together. Thus, multiply 6 by (6/6) to get 36/6, then add (36/6)+(1/6)=(37/6). For his brother's height, you'll want to turn the 5 into a fraction with the same denominator as 3/4, then add them together. Thus, multiply 5 by (4/4) to get 20/4, then add (20/4)+(3/4)=(23/4). Second, you will want to make both fractions have the same denominator. To do this, find the common factor between both of them. The closest one for both 6 and 4 will be 12. Thus, we will be multiplying (37/6) by (2/2) and (23/4) by (3/3) so they both have a denominator of 12, like so: (37/6)*(2/2)=(74/12) (23/4)*(3/3)=(69/12) Now, find the difference! (74/12)-(69/12)=(5/12) There is a 5/12 foot difference between the brothers!